### v1

# Toward a generalization of Lehmer’s problem to adelic curves

Abstract.
In this short note, we investigate the generalization of Lehmer’s problem to finitely generated fields over $\mathbb{Q}$.
Keywords: Lehmer’s problem; adelic curves; heights
MSC: 11G50, 14G40.
1. Introduction
Moriwaki [12] has developed an Arakelov height theory for varieties over a finitely generated extension of a number field. He applied his theory to the study of Bogomolov problem over finitely generated fields over $\mathbb{Q}$ and obtained generalizations of results due to Ullmo [13] and Zhang [14]. As a corollary, he covered the original Raynaud’s theorem (Manin-Mumford’s conjecture). This Arakelov height theory is viewed as an example of an arithmetic intersection theory over an adelic curve in the sense of Chen and Moriwaki [7]. Burgos, Philippon and Sombra [6] have showed that the height of a variety over a finitely generated extension of $\mathbb{Q}$ can be written as an integral of local heights over the set of places of the field. Using their previous work on toric varieties in [5], they gave a combinatorial formula for the height of toric varieties over finitely generated fields over $\mathbb{Q}$.
This work is an attempt to generalize Lehmer’s problem to adelic curves. We show that Lehmer’s problem generalizes to a natural adelic curve defined on the field of rational functions $\mathbb{Q}(T_{1},\ldots,T_{d})$.
Let $f$ be a nonzero polynomial with complex coefficients such that $f(x)=a\prod_{i}(x-\alpha_{i})$. The Mahler measure of $f$ is defined as
(1) $m(f)=\log|a|+\sum_{i}\log\max(1,|\alpha_{i}|).$
If $f$ has integer coefficients then (1) shows that $m(f)$ is nonnegative. A classical theorem of Kronecker asserts that $m(f)=0$ if and only if $|a|=1$ and $\alpha_{i}$ are zero or roots of the unity.
Lehmer asked whether there exists a constant $c>1$ such that for every polynomial $f\in\mathbb{Z}[x]$ with $m(f)>0$, then $m(f)\geq c$. Lehmer’s question has not yet been settled. The best result in the direction of solving Lehmer’s problem is Dobrowolski’s result [10] which gives a lower bound depending on the degree of $f$.
Let $h:\overline{\mathbb{Q}}\rightarrow\mathbb{R}$ be the absolute logarithmic height. Then, Lehmer’s problem can be reformulated as follows. Let $\alpha$ be an algebraic number of degree $d$ over $\mathbb{Q}$ and suppose that it is not a root of unity, then there exists an absolute constant $c>0$ such that
$h(\alpha)\geq\frac{c}{d}.$
Analogues of this conjecture have considered in other contexts, see [9, 8, 11].
Let us point out that generalizations of Lehmer’s problem to higher dimensional varieties have been investigated by Amoroso and David, see for instance [1, 2, 3].
Chen and Moriwaki [7] have introduced the theory of adelic curves. For example, a number field can be viewed as an example of adelic curves. More interestingly, they developed an arithmetic intersection theory on these curves. For instance, one can define a height function on the set of points of adelic curves.
In this paper, we show that Lehmer’s problem generalizes to a natural adelic curve defined on the field of rational functions with coefficients in $\mathbb{Q}$. Let us denote by $S$ the adelic curve on $\mathbb{Q}(T_{1},\ldots,T_{d})$ defined in Section 2.2. In Theorem 4.1, we establish a generalization of Kronecker’s theorem to $S$.
$\{\alpha\in(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
where $h_{S}(\alpha)$ denotes is the height of the point. $[1:\alpha]$ in $\mathbb{P}^{1}(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})$, see Paragraph 2.1 or [12, Paragraph 3.2]. $h_{S}$ is called naive height function in [12].
Assuming Lehmer’s conjecture, then for every $\alpha$ in an algebraic closure of $\mathbb{Q}(T_{1},\ldots,T_{d})$ which is not a root of unity,
$h_{S}(\alpha)\geq\frac{c}{\deg(\alpha)},$
where $h_{S}$ is the standard height function (see (1)), $\deg(\alpha)=[\mathbb{Q}(T_{1},\ldots,T_{d})(\alpha):\mathbb{Q}(T_{1},\ldots,T_{d})]$ and $c$ is an absolute constant, see Theorem 4.3.
2. Adelic curves
In this section, we briefly recall the theory of adelic curves. Let $K$ be a commutative ring and $M_{K}$ be the set of all absolute values on $K$. Following [7], an adelic structure on $K$ consists of a measure space $(\Omega,\mathcal{A},\nu)$ equipped with a map $\phi:\omega\mapsto|\cdot|_{\omega}$ from $\Omega$ to $M_{K}$ satisfying the following properties:
  1. (i)
    $\mathcal{A}$ is a $\sigma$-algebra on $\Omega$ and $\nu$ is a measure on $(\Omega,\mathcal{A})$;
  2. (ii)
    for any $a\in K^{\ast}:=K\setminus\{0\}$, the function $\omega\mapsto\log|a|_{\omega}$ is $\mathcal{A}$-measurable, integrable with respect to the measure $\nu$.
This data is denoted by $S=(K,(\Omega,\mathcal{A},\nu),\varphi)$.
$S$ is said to be proper if the equality
$\int_{\Omega}\log|a|_{\omega}d\nu(\omega)=0,$
holds for each $a\in K^{\ast}$.
From now on, $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ is a proper adelic curve and $K^{ac}$ an algebraic closure of $K$. Let $S\otimes_{K}K^{ac}=({K^{\mathrm{ac}}},(\Omega_{K^{\mathrm{ac}}},\mathcal{A}_{K^{\mathrm{ac}}},\nu_{K^{\mathrm{ac}}}),\phi_{K^{\mathrm{ac}}})$ be the algebraic extension of $S$ by ${K^{\mathrm{ac}}}$ (see [7, Paragraph 3.4.2] for the construction).
2.1. Height of points in $\mathbb{P}^{n}(K^{\mathrm{ac}})$
For $[a_{0}:\ldots:a_{n}]\in\mathbb{P}^{n}({K^{\mathrm{ac}}})$, the height of $[a_{0}:\ldots:a_{n}]$ with respect to the adelic curve $S$ is given as follows.
$h_{S}([a_{0}:\ldots:a_{n}])=\int_{\Omega^{ac}}\log\max(|a_{0}|_{\chi},\ldots,|a_{n}|_{\chi})\nu_{K^{\mathrm{ac}}}(d\chi),$
see [7, Paragraph 3.5] for more details about the construction.
For every $\alpha\in K^{\textrm{ac}}$, we set
(1) $h_{S}(\alpha):=h_{S}([1:\alpha]).$
Lemma 2.1.
The following set
$\{\alpha\in(K^{\mathrm{ac}})^{\ast}\mid h_{S}(\alpha)=0\},$
is a subgroup of $(K^{\mathrm{ac}})^{\ast}$.
Proof.
This follows from the following inequality
$\max(1,xy)\leq\max(1,x)\max(1,y),\quad\forall x,y\geq 0.$
and the product formula (i.e. the properness of $S$).
2.2. The adelic structure on $\mathbb{Q}(T_{1},\ldots,T_{d})$
Let $K=\mathbb{Q}(T_{1},\ldots,T_{d})$ be the field of rational functions of $d$ variables $T_{1},\ldots,T_{d}$ and with coefficients in $\mathbb{Q}$. $\mathbb{Q}(T_{1},\ldots,T_{d})$ can be endowed with an interesting adelic structure, see [7, Paragraph 3.2.6], or [12, Section 3] and [6, Example 2.5] for a more explicit construction. Let us review in details the case when $d=1$.
Let $\Omega=\Omega_{h}\coprod\mathcal{P}\coprod[0,1]_{\ast}$, where $\Omega_{h}$ is the set of all closed points of $\mathbb{P}_{\mathbb{Q}}^{1}\setminus\{\infty\}$, and $\mathcal{P}$ denotes the set all prime numbers, and $[0,1]_{\ast}$ the set of $t\in[0,1]$ such that $e^{2\pi it}$ is transcendental.
Let $\phi:\Omega\rightarrow M_{K}$ be the map sending $\omega\in\Omega$ to $|\cdot|_{\omega}$, defined as follows. Let $x\in\Omega_{h}$. $x$ defines a discrete valuation on $K$ which is denoted by $\mathrm{ord}_{x}(\cdot)$. Let $F_{x}$ be an irreducible polynomial in $\mathbb{Z}[T]$ with $F_{x}(x)=0$. Let $H(x)$ be the Mahler measure of the polynomial $F_{x}$. It is defined as follows.
$H(x):=\exp\left(\int_{0}^{1}\log|F_{x}(e^{2\pi it})|dt\right).$
Let $|\cdot|_{x}$ be the absolute value on $\mathbb{Q}(T)$ such that
$|\varphi|_{x}=H(x)^{-\mathrm{ord}_{x}(\varphi)},\quad\forall\varphi\in\mathbb{Q}(T).$
For $p$ a prime number, let $|\cdot|_{p}$ be $p$-adic value on $\mathbb{Q}$. It extends to $\mathbb{Q}[T]$ as follows.
$|\varphi|_{p}:=\max_{j=0,\ldots,d}|a_{j}|_{p},\quad\forall\varphi=\sum_{i=0}^{d}a_{i}T^{i}\in\mathbb{Q}[T],$
and then, by multiplicativity, to $\mathbb{Q}(T)$.
For any $t\in[0,1]_{\ast}$, let $|\cdot|_{t}$ be the absolute value on $\mathbb{Q}(T)$ such that
$|\varphi|_{t}:=|\varphi(e^{2\pi it})|,\quad\forall\varphi\in\mathbb{Q}(T),$
where ||$\cdot$|| ⋅ | denotes the usual absolute value of $\mathbb{C}$.
$\Omega_{h}$ and $\mathcal{P}$ are equipped with the discrete $\sigma$-algebras, and $[0,1]_{\ast}$ with the restriction of the Borel $\sigma$-algebra on $[0,1]$. Let $\mathcal{A}$ be the $\sigma$-algebra on $\Omega$ generated by the above $\sigma$-algebras on $\Omega_{h},\mathcal{P}$ and $[0,1]_{\ast}$ respectively. Let $\nu$ be the measure on $\Omega$ such that $\nu(\{x\})=1$ for $x\in\Omega_{h}$, that $\nu(\{p\})=1$ for any prime number $p$ and that the restriction of $\nu$ on $[0,1]_{\ast}$ coincides with the Lebesgue measure.
This data defines an adelic structure on $\mathbb{Q}(T)$. Moreover, $(\mathbb{Q}(T),(\Omega,\mathcal{A},\nu),\phi)$ is proper, see [7, Paragraph 3.2.5].
3. On Lehmer’s problem on adelic curves
Question 3.1.
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve and $K^{ac}$ an algebraic closure of $K$. Is there a constant $c>0$ such for every $\alpha\in K^{\mathrm{ac}}$ with $h_{S}(\alpha)>0$, then
$h_{S}(\alpha)\geq\frac{c}{[K(\alpha):K]}.$
When $K$ is $\mathbb{Q}$ and $S$ is the classical adelic curve associated with $\mathbb{Q}$, then this is the classical Lehmer’s problem.
Next, we show that the classical Lehmer’s problem generalizes naturally to the adelic curve $\mathbb{Q}(T)$.
4. Lehmer’s problem on $\mathbb{Q}(T_{1},\ldots,T_{d})$
Let $S$ be the adelic curve on $\mathbb{Q}(T)$ recalled in Paragraph 2.2. Let
$[\varphi:\psi]\in\mathbb{P}^{1}\left(\overline{\mathbb{Q}(T)}\right).$
We may assume that $\varphi$ and $\psi$ are integral over $\mathbb{Z}[T]$. Then, there exist $d,e\in\mathbb{N}$ and $f_{1}(T),\ldots f_{d}(T)$ and $g_{1}(T),\ldots,g_{e}(T)$ in $\mathbb{Z}[T]$ such that
$\varphi^{d}+f_{1}(T)\varphi^{d-1}+\ldots+f_{d-1}(T)\varphi+f_{d}(T)=0,$
and
$\psi^{e}+g_{1}(T)\varphi^{e-1}+\ldots+g_{e-1}(T)\psi+g_{e}(T)=0.$
Let
$c_{\varphi,\psi}:=\gcd(c(f_{d})^{e},c(g_{e})^{d}).$
where $c(f_{d})$ (resp. $c(g_{e})$) is the common divisor of the coefficients of $f_{d}$ (resp. $g_{e}$).
If $\{x\in\overline{\mathbb{Q}}\mid g_{e}(x)=f_{d}(x)=0\}$ is not empty, we let
$F(T):=\prod_{\begin{subarray}{c}x\in\overline{\mathbb{Q}}\\ g_{e}(x)=f_{d}(x)=0\end{subarray}}F_{x}(T)^{\min(\mathrm{ord}_{x}(f_{d}^{e}),\mathrm{ord}_{x}(g_{e}^{d}))}\,,$
with $F_{x}\in\mathbb{Z}[T]$ with coprime coefficients, irreducible in $\mathbb{Q}[T]$ and $F_{x}(x)=0$ for $x\in\overline{\mathbb{Q}}$. Otherwise, we let
$F(T)=1.$
By the product formula, we have
$h_{S}([\varphi:\psi])=h_{S}\Bigl{(}\Bigl{[}\tfrac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}:\tfrac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{]}\Bigr{)}.$
Let $K^{\prime}$ be a finite extension of $\mathbb{Q}(T)$ such that $\varphi$ and $\psi$ are in $K^{\prime}$.
Let $\omega\in\Omega_{K^{\prime}}$ with $\omega\mid x$ where $x\in\Omega_{h}$. If $x$ is a common zero of $f_{e}$ and $g_{d}$, then
$\begin{split}\log\left|\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}=&\frac{1}{d}\log|f_{d}|_{x}-\frac{1}{de}\log|F|_{x}\\ =&-\frac{\mathrm{ord}_{x}(f_{d}(T))}{d}\log H(x)+\frac{\min(\mathrm{ord}_{x}(f_{d}(T)^{e}),\mathrm{ord}_{x}(g_{e}(T)^{d}))}{de}\log H(x).\end{split}$
and
$\begin{split}\log\left|\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{d}}}\right|_{\omega}=&\frac{1}{e}\log|g_{e}|_{x}-\frac{1}{de}\log|F|_{x}\\ =&-\frac{\mathrm{ord}_{x}(g_{e}(T))}{e}\log H(x)+\frac{\min(\mathrm{ord}_{x}(f_{d}(T)^{e}),\mathrm{ord}_{x}(g_{e}(T)^{d}))}{de}\log H(x).\end{split}$
We conclude that
$\max\left(\left|\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega},\left|\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}\right)=1$
This equation holds if $x$ is not a common zero of $f_{e}$ and $g_{d}$.
Let $\omega\in\Omega$ with $\omega\mid p$ and $p\in\mathcal{P}$. We have
$\left|\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}=\frac{|f_{d}(T)|_{p}^{\frac{1}{d}}}{|(c_{\varphi,\psi})^{\frac{1}{de}}|_{p}}=\left(\frac{|c(f_{d})^{e}|_{p}}{|c_{\varphi,\psi}|_{p}}\right)^{\frac{1}{de}},$
and
$\left|\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}=\frac{|g_{e}(T)|_{p}^{\frac{1}{e}}}{|(c_{\varphi,\psi})^{\frac{1}{de}}|_{p}}=\left(\frac{|c(g_{e})^{d}|_{p}}{|c_{\varphi,\psi}|_{p}}\right)^{\frac{1}{de}}.$
We conclude that
$\max\left(\left|\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega},\left|\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right|_{\omega}\right)=1.$
Case of $w\in\Omega$ with $w\mid t$ and $t\in[0,1]_{\ast}$. Note that $K^{\prime}_{\omega}=\mathbb{Q}(T)_{t}=\mathbb{C}$. We have then
$\begin{split}\int_{\begin{subarray}{c}w\in\Omega\\ w\mid t,\ t\in[0,1]_{\ast}\end{subarray}}\log\max\Bigl{(}&\Bigl{|}\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega},\Bigl{|}\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega}\Bigr{)}d\nu(\omega)\\ =&\int_{t\in[0,1]_{\ast}}\frac{1}{[K^{\prime}:\mathbb{Q}(T)]}\left(\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\max\Bigl{(}\Bigl{|}\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega},\Bigl{|}\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega}\Bigr{)}\right)dt.\end{split}$
Gathering the above computations, we obtain that
$\frac{1}{[K^{\prime}:\mathbb{Q}(T)]}\max\Bigl{(}\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\Bigl{|}\frac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega}dt,\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\Bigl{|}\frac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\Bigr{|}_{\omega}dt\Bigr{)}\leq h_{S}([\varphi:\psi]).$
Using [7, Lemma 3.5.4], we obtain
$\begin{split}\tfrac{1}{[K^{\prime}:\mathbb{Q}(T)]}\max\Bigl{(}\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}\log\Bigl{|}\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr{|}_{\omega}dt,\int_{t\in[0,1]_{\ast}}\sum_{\begin{subarray}{c}w\in K^{\prime}\\ w\mid t\end{subarray}}&\log\Bigl{|}\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr{|}_{\omega}dt\Bigr{)}\\ &\leq[K^{\prime}:\mathbb{Q}(T)]h_{S}([\varphi:\psi]).\end{split}$
Hence
(2) $\begin{split}\max\Bigl{(}\int_{t\in[0,1]}\log\Bigl{|}\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\varphi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr{|}_{t}dt,\int_{t\in[0,1]}&\log\Bigl{|}\mathrm{N}_{K^{\prime}/\mathbb{Q}(T)}\left(\tfrac{\psi}{(c_{\varphi,\psi}F)^{\frac{1}{de}}}\right)\Bigr{|}_{t}dt\Bigr{)}\\ &\leq[K^{\prime}:\mathbb{Q}(T)]h_{S}([\varphi:\psi]).\end{split}$
Theorem 4.1.
(A Kronecker’s theorem)
$\{\alpha\in(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
Proof.
For simplicity, we prove the theorem when $d=1$. Let $\varphi$ and $\psi$ be two coprime polynomials in $\mathbb{Z}[T]$. We have
$h_{S}([\varphi:\psi])=\int_{0}^{1}\log\max(|\varphi(e^{2\pi it})|,|\psi(e^{2\pi it})|)dt.$
Let us assume that $h_{S}([\varphi:\psi])=0$. This implies that
$\varphi(z),\psi(z),\ \text{and}\ \varphi(z)+y\psi(z)$
are generalized cyclotomic polynomials in the variables $y,z$ (see for instance [4]). We have used the following fact:
$\int_{0}^{1}\log\max(|\varphi(e^{2\pi it})|,|\psi(e^{2\pi it})|)dt=\int_{0}^{1}\int_{0}^{1}\log|\varphi(e^{2\pi it})+e^{2\pi is}\psi(e^{2\pi it})|dtds.$
It is not difficult to show that $\phi$ and $\psi$ are, in fact, roots of unity in $\overline{\mathbb{Q}}$.
Let $\alpha\in\overline{\mathbb{Q}(T)}$ with $h_{S}(\alpha)=0$. Using (2), we can conclude that $\alpha$ is a root of unity in $\overline{\mathbb{Q}}$.
Let $\varphi\in\overline{\mathbb{Q}(T)}$. Let $d:=[\mathbb{Q}(T)(\varphi):\mathbb{Q}(T)]$. We can assume
$[1:\varphi]=[\varphi_{1}:\varphi_{2}],$
with $\varphi_{1}\in\mathbb{Z}[T]$ and $\varphi_{2}\in\overline{\mathbb{Z}[T]}$ (integral closure of $\mathbb{Z}[T]$).
From (2) and assuming that Lehmer’s problem is true, we get
$h_{S}(\varphi)(=h_{S}([\varphi_{1}:\varphi_{2}]))\geq\frac{c}{\deg(\varphi)},$
where $\deg\varphi=[\mathbb{Q}(T)(\alpha):\mathbb{Q}(T)]$ and $c$ is the constant in Lehmer’s problem.
Remark 4.2.
If $\alpha\in\overline{\mathbb{Q}}$, then
$h_{S}(\alpha)=h(\alpha),$
where $h$ is the standard Weil’s height of $\alpha$.
We conclude that
Theorem 4.3.
Lehmer’s problem is equivalent to Lehmer’s problem on $\mathbb{Q}(T_{1},\ldots,T_{d})$.